# MAVias: Mitigate any Visual Bias

### v1

@@ 4.3 Ablation Analysis
w/o regularizer 89.4$\pm$1.6 95.2$\pm$0.4
w/ regularizer 93.7$\pm$0.4 94.5$\pm$0.4
Note that in the Appendix, we also investigate different methods for computing embeddings from tags and examine the effect of the tags’ vocabulary size.
5 Conclusion
In this work, we presented a framework to address open-set biases in CV using instance-level descriptive tags. Unlike traditional methods relying on predefined biases, MAVias offers a flexible, scalable solution for identifying and mitigating unknown biases. It detects visual features that may introduce bias and reduces their impact through a robust approach. Extensive experiments show MAVias outperforms existing methods in detecting and mitigating complex, real-world biases. The main requirement is an image tagging model with a comprehensive tag vocabulary and an LLM suited to the data context. For example, cloud-deployed models are not ideal for sensitive data like medical images.
Appendix A Appendix
A.1 Theoretical Justification
A.1.1 Impact on Gradients
where $\sigma^{(k)}$ denotes the $k$th class’ softmax score.
Combining Eq. 4 and Eq. 5 we get:
where
which indicates that through our framework the main model $f_{\boldsymbol{\theta}}$ is enforced to update its weights based mainly on the gradients corresponding to the bias-conflicting samples $\mathbf{x}^{(c)}$. This essentially means that $f_{\boldsymbol{\theta}}$ will eventually ignore the visual cues prevalent in $\mathbf{x}^{(a)}$ samples corresponding to irrelevant information encoded by the tag model $g_{\boldsymbol{\phi}}$.
A.1.2 Unbiased Distribution Learning
Let $p_{tr}(X,Y,B)$ be the training data distribution where $X$, $Y$, $B$ are random variables associated with the data, the label and the bias attribute, respectively. From Theorem 1 of [16] we know that if $P_{u}$ is a shift of $p_{tr}$ such that the distribution $P_{u}(Y|B)$ is uniform (and therefore unbiased), then for a model $f_{\boldsymbol{\theta}}=p_{u}(y|\mathbf{x})$ that estimates the conditional probability over $P_{u}$, its probability over $p_{tr}$ will be
In the case of MAVias, let the probability $p_{tr}(y|b)$ be estimated through the projection and classification layers, $\mathbf{z}_{\text{tag}}=f_{\boldsymbol{\theta}_{c}}(g_{\boldsymbol{\phi}}(\mathbf{e}))$ which use the visual bias embeddings $\mathbf{e}$ (encoding the irrelevant tags) to produce logits $\mathbf{z}_{\text{tag}}$. As a result,
where $A(\mathbf{z}_{\text{tag}})=\sum_{y^{\prime\prime}}e^{\mathbf{z}_{\text{tag}}(y^{\prime\prime})}$. Replacing Eq. (9) to Eq. (8) gives
thus leading to the logit addition of Eq. (2).
For this approach to work, however, one must ensure that models $g_{\boldsymbol{\phi}}$ and $f_{\boldsymbol{\theta}_{c}}$ are successfully trained to produce logits $\mathbf{z}_{\text{tag}}$ that predict $p_{tr}(y|b)$ and are not dominated by $\mathbf{z}_{\text{main}}$. This is achieved by the regularization term of Eq. (3) which controls the magnitude differences of $\mathbf{z}_{\text{main}}$ and $\mathbf{z}_{\text{tag}}$.
A.2 Experimental Setup
A.2.1 Model architectures
For comparability purposes, we employ the same network architectures as those used in other works [39, 16, 2, 40]. In particular, for CelebA, we adopt the ResNet-18 architecture [14], for Waterbirds, UrbanCars, and ImageNet9 datasets we use Resnet-50 networks. In all experiments, the projection layer is a dense layer that gets vision-language embeddings as input and its output size is aligned to the feature size of the main model.
A.2.2 Implementation details
The SGD optimizer is employed for all datasets except for CelebA, where Adam optimizer is used. We use an initial learning rate of 0.001, which is divided by 10 every 1/3 of the training epochs. The weight decay is set to $10^{-4}$. The batch size is 128 for CelebA and 64 for Waterbirds, UrbanCars, and ImageNet9. Following previous works [16, 37, 24], we train the models for 40, 100, 300, and 40 for CelebA, Waterbirds, UrbanCars, and ImageNet9, respectively, For Waterbirds and UrbanCars, we do not use a learning rate scheduler. The parameters $(\alpha,\lambda)$ are (0.01, 0.5), (0.05, 0.6), (0.01, 0.4), and (0.001, 0.7) for CelebA, Waterbirds, UrbanCars, and ImageNet9, respectively. The optimal values of the hyperparameters are derived through a grid search. RAM [49], ChatGPT-4o [33], CLIP [36] are employed for image tagging, irrelevant tag filtering, and vision-language encoding, respectively. Experiments on CelebA, Waterbirds, and UrbanCars, were conducted on an NVIDIA A100 GPU, while experiments on Imagenet were conducted on a single NVIDIA RTX-4090 GPU. All experiments were repeated for 5 different random seeds.
A.2.3 Prompting
System Prompt I will provide you with the name of a target class and a large list of tags. Your task is to evaluate the tags and identify only those directly related to the target class. A tag is considered relevant if it describes or is an essential part of the object associated with the class name. This includes tags that refer to: physical components, defining features, inherent characteristics, and essential behaviors or functions of the object. For example, if the target class is “bee” tags like “insect”, “wing”, and “buzz” are relevant because they describe core aspects of what a bee is or does. Conversely, a tag is irrelevant if it refers to elements that are not an intrinsic part of the object. Irrelevant tags may include: background details, environmental context, colors (unless a defining characteristic), lighting, textures, other objects, or other non-essential contextual elements. For example, in the case of the class “bee”, tags like “sky”, “flower”, or “blue” would be irrelevant, as they describe the environment or background rather than the bee itself. Please output only the relevant tags in JSON format only (i.e., { relevant_tags: [the list of tags]}).
Figure 5: LLM system prompt for deriving the relevant tags.
A.3 Two-moon Distribution Experiment
(a) Vanilla predictions on axes $x$ and $y$.
(b) Vanilla predictions on axes $x$ and $z$.
(c) MAVias predictions on axes $x$ and $y$.
(d) MAVias predictions on axes $x$ and $z$.
A.4 Comparative Analysis
This section provides a comprehensive comparative analysis involving both bias label-aware (BA) and bias label-unaware (BU) methodologies. As shown in Tab. 10, MAVias demonstrates competitive performance on CelebA compared to the best-performing BA method, i.e., -0.6% bias-conflict accuracy and -1.5% unbiased accuracy, even though MAVias has no access to bias-related information.
Table 10: Evaluation of the proposed method on CelebA with BlondHair and gender as biased attribute and target, respectively. Bold denotes the best-performing bias label-unaware (BU) method and underlined denotes the best-performing bias label-aware method.
Methods BU Bias
BlondHair
Unbiased Bias-conflict
LNL [19] 80.1$\pm$0.8 61.2$\pm$1.5
DI [45] 90.9$\pm$0.3 86.3$\pm$0.4
EnD [42] 86.9$\pm$1.0 76.4$\pm$1.9
BC-BB [16] 91.4$\pm$0.0 87.2$\pm$0.2
FairKL [2] 81.7$\pm$1.7 69.9$\pm$2.4
FLAC [39] 91.2$\pm$0.3 88.7$\pm$0.5
LfF [31] 84.2$\pm$0.3 81.2$\pm$1.4
SoftCon [16] 84.1 74.4
FLAC-B [39] 87.0$\pm$0.6 84.9$\pm$2.2
MAVias 89.7$\pm$0.6 87.1$\pm$1.7
Furthermore, Tab. 11 evaluates MAVias on Waterbirds, a well-known benchmark for spurious correlation problems between bird type and background environment. Notably, MAVias attains 93.7% Worst-Group accuracy, outperforming all compared BA methods, i.e., GroupDRO, BAdd, and DFR. Also, MAVias achieves this improvement while maintaining competitive average accuracy (+0.3%), which indicates that the reduction in bias did not come at the cost of overall performance. In contrast, BU methods like JTT and DISC, which can be directly compared with MAVias, show a larger gap in WG performance, indicating their limited effectiveness as a result of the lack of bias information (i.e., type of bias or bias labels), a limitation that MAVias successfully addresses.
Table 11: Evaluation on Waterbirds.
Method BU WG Acc. Avg. Acc.
GroupDro [37] 90.6$\pm$1.1 91.8$\pm$0.3
BAdd [40] 92.9$\pm$0.3 93.6$\pm$0.2
DFR [35] 92.9$\pm$0.2 94.2$\pm$0.4
JTT [26] 86.7$\pm$1.5 93.3$\pm$0.3
DISC [46] 88.7$\pm$0.4 93.8$\pm$0.7
MAVias 93.7$\pm$0.4 94.5$\pm$0.4
Finally, Tab. 12 presents the performance of the relevant methods on the UrbanCars dataset. This benchmark poses a greater challenge due to the presence of multiple biases, as reflected in the results, i.e., most BU methods perform significantly worse than BA methods. However, MAVias achieves competitive performance, with only 2%, 0.8%, and 2.8% drops in BG, CoObj, and BG+CoObj Gaps, respectively, compared to the best-performing BA methods.
Table 12: Evaluation on UrbanCars.
Method BU I.D. Acc BG Gap CoObj Gap BG+CoObj Gap
GroupDro [37] 91.6 -10.9 -3.6 -16.4
DFR [35] 89.7 -10.7 -6.9 -45.2
LLE [24] 96.7 -2.1 -2.7 -5.9
BAdd [40] 91.0$\pm$0.7 -4.3$\pm$0.4 -1.6$\pm$1.0 -3.9$\pm$0.4
LfF [31] 97.2 -11.6 -18.4 -63.2
JTT [26] 95.9 -8.1 -13.3 -40.1
Debian [23] 98.0 -14.9 -10.5 -69.0
MAVias 92.8$\pm$0.8 -4.1$\pm$0.6 -2.4$\pm$1.4 -6.7$\pm$1.4
A.5 Ablation Study
First, let us report the accuracy of $g_{\boldsymbol{\phi}}(\cdot)$ on the training samples of the urban car class from the UrbanCars dataset, which correlates the target classes with relevant backgrounds and co-occurring objects. As shown in Tab. 13, the $g_{\boldsymbol{\phi}}$ shows a clear distinction in its performance: its accuracy increases from 0% for bias-conflicting samples (i.e., those with a country background and co-occurring object) to 98.42% for bias-aligned samples (i.e., those with an urban background and co-occurring object). This result confirms that our tag-based approach produces embeddings that effectively capture bias, providing a foundation for the main model $f_{\boldsymbol{\theta}}(\cdot)$ to focus on learning unbiased features.
Table 13: MAVias: $g_{\boldsymbol{\theta}}(\cdot)$ accuracy for the UrbanCars training samples belonging to subgroups of urban car class with different biases.
@@ A.5 Ablation Study
Furthermore, we investigate the impact of the vocabulary size employed for image tagging. The original vocabulary RAM uses has a size of 4585 words. Tab. 14 reports the performance of MAVias for different portions of the original vocabulary. Notably, using 30% or more of the original vocabulary yields highly effective models, which is due to the dataset’s bias (aquatic or terrestrial environments) that a variety of tags can capture.
@@
Moreover, Tab. 15 reports the performance of MAVias when computing individual embeddings for each tag and then averaging them compared to the one computing a single embedding for all tags combined. As one may observe, although both are highly effective, the latter can provide more representative features.

### v2

@@ 4.6 Ablation Analysis
[…]
In addition, let us report the accuracy of $g_{\boldsymbol{\phi}}(\cdot)$ on the training samples of the urban car class from the UrbanCars dataset, which correlates the target classes with relevant backgrounds and co-occurring objects. As shown in Tab. 11, the $g_{\boldsymbol{\phi}}$ shows a clear distinction in its performance: its accuracy increases from 0% for bias-conflicting samples (i.e., those with a country background and co-occurring object) to 98.42% for bias-aligned samples (i.e., those with an urban background and co-occurring object). This result confirms that our tag-based approach produces embeddings that effectively capture bias, providing a foundation for the main model $f_{\boldsymbol{\theta}}(\cdot)$ to focus on learning unbiased features.
Table 11: MAVias: $g_{\boldsymbol{\theta}}(\cdot)$ accuracy for UrbanCars training samples belonging to subgroups of urban car class with different biases.
@@
Furthermore, we investigate the impact of the vocabulary size employed for image tagging. The original vocabulary of RAM has a size of 4585 words. Tab. 12 reports the performance of MAVias for different portions of the original vocabulary. Notably, using 30% or more of the original vocabulary yields highly effective models.
@@
5 Conclusion
We presented a method to address open-set biases in CV using instance-level descriptive tags. Unlike previous methods relying on predefined biases, MAVias offers a flexible solution for identifying and mitigating unknown biases. It detects visual features that may introduce bias and reduces their impact. Extensive experiments show MAVias outperforms existing methods in detecting and mitigating complex, real-world biases. The main requirement is an image tagging model with a comprehensive tag vocabulary and an LLM suited to the application context. For example, cloud-deployed models are not ideal for applications where sensitive data is involved, e.g., medical images, images of personal identity documents, etc.
Acknowledgments
This research was supported by the EU Horizon Europe projects MAMMOth (grant no. 101070285), ELIAS (grant no. 101120237), and ELLIOT (grant no. 101214398).
[…]
Supplementary Material
1 Theoretical Justification
1.1 Impact on Gradients
where $\sigma^{(k)}$ denotes the $k$th class’ softmax score.
Combining Eq. 3 and Eq. 4 we get:
where
which indicates that through our framework, the main model $f_{\boldsymbol{\theta}}$ is enforced to update its weights based mainly on the gradients corresponding to the bias-conflicting samples $\mathbf{x}^{(c)}$. This essentially means that $f_{\boldsymbol{\theta}}$ will eventually ignore the visual cues prevalent in $\mathbf{x}^{(a)}$ samples corresponding to irrelevant information encoded by the tag model $g_{\boldsymbol{\phi}}$.
1.2 Unbiased Distribution Learning
Let $p_{tr}(X,Y,B)$ be the training data distribution where $X$, $Y$, $B$ are random variables associated with the data, the label, and the bias attribute, respectively. From Theorem 1 of [hong2021bb] we know that if $P_{u}$ is a shift of $p_{tr}$ such that the distribution $P_{u}(Y|B)$ is uniform (and therefore unbiased), then for a model $f_{\boldsymbol{\theta}}=p_{u}(y|\mathbf{x})$ that estimates the conditional probability over $P_{u}$, its probability over $p_{tr}$ will be
In the case of MAVias, let the probability $p_{tr}(y|b)$ be estimated through the projection and classification layers, $\mathbf{z}_{\text{tag}}=f_{\boldsymbol{\theta}_{c}}(g_{\boldsymbol{\phi}}(\mathbf{e}))$, which use the visual bias embeddings $\mathbf{e}$ (encoding the irrelevant tags) to produce logits $\mathbf{z}_{\text{tag}}$. As a result,
where $A(\mathbf{z}_{\text{tag}})=\sum_{y^{\prime\prime}}e^{\mathbf{z}_{\text{tag}}(y^{\prime\prime})}$. Replacing (8) to (7) gives
thus leading to the logit addition of Eq. (1).
For this approach to work, however, one must ensure that models $g_{\boldsymbol{\phi}}$ and $f_{\boldsymbol{\theta}_{c}}$ are successfully trained to produce logits $\mathbf{z}_{\text{tag}}$ that predict $p_{tr}(y|b)$ and are not dominated by $\mathbf{z}_{\text{main}}$. This is achieved by the logit alignment term of Eq. (2) which controls the magnitude differences of $\mathbf{z}_{\text{main}}$ and $\mathbf{z}_{\text{tag}}$.
2 Experimental Setup
2.1 Prompting
System Prompt I will provide you with the name of a target class and a large list of tags. Your task is to evaluate the tags and identify only those directly related to the target class. A tag is considered relevant if it describes or is an essential part of the object associated with the class name. This includes tags that refer to: physical components, defining features, inherent characteristics, and essential behaviors or functions of the object. For example, if the target class is “bee”, tags like “insect”, “wing”, and “buzz” are relevant because they describe core aspects of what a bee is or does. Conversely, a tag is irrelevant if it refers to elements that are not an intrinsic part of the object. Irrelevant tags may include: background details, environmental context, colors (unless a defining characteristic), lighting, textures, other objects, or other non-essential contextual elements. For example, in the case of the class “bee”, tags like “sky”, “flower”, or “blue” would be irrelevant, as they describe the environment or background rather than the bee itself. Please output only the relevant tags in JSON format only (i.e., { relevant_tags: [the list of tags]}).
Figure 4: LLM system prompt for deriving the relevant tags.
3 Two-moon Distribution Experiment
(a) Vanilla outputs on axes $x$ and $y$.
(b) Vanilla outputs on axes $x$ and $z$.
(c) MAVias outputs on axes $x$ and $y$.
(d) MAVias outputs on axes $x$ and $z$.
4 Comparative Analysis
In this section, we compare the performance of MAVias with other competitive methods on closed-set (i.e., predefined bias) scenarios. The reported results encompass both bias-label aware (BA) and bias-label unaware (BU) methods. However, it should be noted that MAVias can be directly compared only with the BU approaches. Table 13 presents the performance of MAVias on CelebA with gender as target and the BlondHair as the biased attribute. The performance on unbiased and bias-conflicting samples indicates the ability of MAVias to reduce reliance on spurious correlations between the target and the corresponding biased attributes. Specifically, MAVias achieves 87.1% accuracy on bias-conflicting samples, surpassing state-of-the-art BU methods such as FLAC-B [sarridis2023flac] and LfF [nam2020LfF], while maintaining high unbiased accuracy (89.7%). Table 14 evaluates MAVias on Waterbirds, a well-known benchmark for spurious correlation problems between bird type and background environment. Notably, MAVias attains 93.7% Worst-Group accuracy, outperforming all compared BU methods (+5%), while slightly improving the average accuracy (+0.7%). The results on UrbanCars in Tab. 15 further validate the efficacy of MAVias in mitigating multiple biases introduced by co-occurring objects and background features. While the in-distribution accuracy (I.D. Acc) of MAVias is marginally lower than that of LfF [nam2020LfF] or Debian [li2022discover], it achieves substantial improvements in reducing the BG Gap (4.1) and CoObj Gap (2.4), as well as their combined effects (BG+CoObj Gap: 6.7).
Table 13: Bias-conflict and Unbiased accuracy comparison on CelebA with BlondHair and gender as biased attribute and target, respectively. Bold denotes the best-performing bias label-unaware (BU) method and underlined denotes the best-performing bias label-aware method.
Methods BU Bias
BlondHair
Unbiased Bias-conflict
[…]
SoftCon [hong2021bb] 84.1 74.4
[…]
Table 14: Evaluation on Waterbirds.
Method BU WG Acc. Avg. Acc.
[…]
Table 15: Evaluation on UrbanCars.
Method BU I.D. Acc BG Gap CoObj Gap BG+CoObj Gap
GroupDro [sagawa2019distributionally] 91.6 10.9 3.6 16.4
DFR [qiu2023simple] 89.7 10.7 6.9 45.2
LLE [li2023whac] 96.7 2.1 2.7 5.9
[…]
LfF [nam2020LfF] 97.2 11.6 18.4 63.2
JTT [liu2021just] 95.9 8.1 13.3 40.1
Debian [li2022discover] 98.0 14.9 10.5 69.0
[…]
Figure 6: Impact of hyperparameters $\alpha$ and $\lambda$ on the worst-group accuracy (closed-set protocol) on the Waterbirds dataset.
5 Ablations & Hyperparameter Tuning
@@
Tab. 16 reports the performance of MAVias when computing individual embeddings for each tag and then averaging them compared to the one computing a single embedding for all tags combined. As one may observe, although both are highly effective, the latter can provide more representative features.